# Stability of dark solitons in a bubble Bose-Einstein condensate

### v1

@@ Abstract
The dynamic stability of dark solitons trapped on the surface of a two-dimensional spherical bubble is investigated. In this spherical geometry of the Bose-Einstein condensate, dark solitons are found to be unstable for the interaction parameter $\varepsilon\gtrsim 8.37$, since discrete angular modes drive snake instabilities, with the generation of vortex dipoles. We show analytically and numerically that, for each angular mode $m\geq 2$, there exists exactly one unstable mode whose dominance determines the number $m$ of vortex dipoles. Time-dependent simulations confirm the formation of vortex dipoles.
@@
Bose-Einstein condensed gases in spherical geometries have recently attracted attention due to experiments performed aboard the International Space Station in a microgravity environment with ultra-cold gases confined in spherical and/or ellipsoidal surfaces [1, 2]. These experimental investigations were designed based on previous theoretical work on shell-like potentials [3, 4, 5]. Ground-based experiments also revealed shell bubbles, by exploring two species [6], or intending to observe alternative two-dimensional (2D) closed geometries [7]. Such investigations have shed new light on the physics of low-dimensional quantum gases, especially in closed 2D shells. See Ref. [8] for a recent overview of the present status and perspectives on quantum gases in bubble traps. Several interesting problems have been studied in this context, considering fundamental physics properties in shell-like structures [9, 10, 11, 12, 13, 14], vortex dynamics and stability [16, 15, 17], dipole interaction effects [18, 19], Berezinskii-Kosterlitz-Thouless (BKT) transition [20], thermodynamic properties of the gas adiabatic expansion from filled sphere to hollow one [21], and to our main interest, some attention has been paid to the properties of condensate mixtures trapped on a bubble [23, 22, 24]. Furthermore, the study of dimers on a spherical surface shows that the dimers are squeezed in the direction orthogonal to the center of mass motion, qualitatively changing its geometry, from 2D to one-dimensional (1D), leading to two-soliton motion on a bubble surface [25].
@@
In the context of the properties of Bose-Einstein condensates (BECs) on spherical surfaces, a topic of particular interest concerns the existence, stability, and propagation of dark solitons [26]. Bubble BECs are realized when atoms are confined to a thin spherical shell, creating unique curvature effects that influence the dynamics of solitons. Dark solitons are shape-localized propagating waves manifested by depressions in the condensate density. On a bubble surface, geometric effects can alter the instability dynamics of dark solitons, as compared to their propagation on flat surfaces. The curved surface is expected to affect their dynamics, propagation speed, and stability. Unlike in flat quasi-1D or quasi-2D BECs, where dark solitons are susceptible to snake instability [27], decaying into vortices, the closed topology of a bubble is expected to alter the decay pathways, with the curvature introducing geometric constraints affecting the stability dynamics. Therefore, the soliton can experience strain due to the curvature, leading to contraction or expansion depending on the interactions.
@@
Dark solitons in a flat BEC are known to oscillate and decay due to quantum fluctuations or dissipation. In quasi-1D traps, dark solitons can be metastable and decay due to transverse instabilities [28, 29]. Using harmonic traps, they were studied in [30, 31], both analytically and computationally, where it was found that the large-amplitude field modulations at a frequency resonant with the energy of a dark soliton give rise to a state with multiple vortices. The stability spectrum of the dark soliton contains complex frequencies, which disappear for sufficiently small numbers of atoms or for a large transverse confinement. The relationship between these complex modes and the snake instability was investigated numerically by real-time propagation, see also  [32]. In ring geometries, dark solitons in BEC systems, introduced in [33], were further investigated in [34].
@@
Two-component BECs were considered in [35], where the dark soliton exists in one of the condensate components and the soliton nodal plane is filled with the second component. The filled solitons are stable for hundreds of milliseconds. By selectively removing the filling, one can make the soliton more susceptible to dynamical instabilities. For a condensate in a spherically symmetric potential, these instabilities cause the dark soliton to decay into stable vortex rings. By studying the oscillations and interactions of dark and dark-bright solitons, it was shown in  [36] that the stability of solitons can be controlled in BECs, via confinement.
@@
The aim of the research we present concerns a study on the spectral stability of dark solitons under discrete $m$-angular modes within a BEC confined on the surface of a bubble in the approximation of a 2D spherical hollow shell [15]. We show the occurrence of exactly one unstable mode for each $m\geq 2$. These unstable modes induce snake-like excitations, which break the condensate into pieces, as observed in [15]. As verified numerically, an $m$-dominated unstable mode is followed by the creation of $m$ vortex-antivortex pairs, in agreement with our analytical predictions.
@@
This letter is organized as follows. We first present the theoretical model for an atomic BEC confined to the surface of a perfect sphere, named a bubble. Dark solitons, together with the asymptotic approximations in the limit of small and large chemical potentials, are obtained analytically and numerically, supported by Supplemental Material (SM) [37]. We then report the analytical results of the corresponding stability analysis, demonstrating they are in excellent agreement with numerical observations. Finally, we report outcomes of the nonlinear dynamics of the system, from which it follows that the instabilities result in the formation of vortex dipoles on the dark soliton condensate.
@@
Mathematical model — The present study is performed by assuming the condensate is trapped on the surface of a rigid spherical shell, aiming to mimic the cold-atom bubble experiments that are currently being performed in microgravity environments. In this approach, the system can be studied by considering a reduction of the 3D Gross-Pitaevskii equation (GPE) to a corresponding 2D system described by spherical coordinates $\theta\in[0,\pi]$ and $\phi\in[0,2\pi]$, with fixed radius $R>0$, which is also assumed to be the space unit. In the dimensional reduction, $\delta R$ is also taken as being the radial thickness of the spherical bubble. The two-dimensional approximation is reasonable since excitations in the radial direction are inaccessible due to the large amount of energy required for them. This is true when the thickness $\delta R$ of the 3D spherical shell is very small compared to its radius $R$, that is, $\delta R\ll R$, as discussed in Refs.[15, 22]. We also stress that our main concern is the dynamic stability of the dark soliton. We do not take into account its energetic instability, which could be triggered by a thermal cloud that is neglected here. The condensate can be described in the mean-field approach as a system of two coupled GPEs [38, 39] with the nonlinear two-body parameter $\overline{g}\equiv{4\pi\hbar^{2}{a}N}/{M}$, where $N$ is the number of atoms with mass $M$, and $a$ is the atom-atom $s$-wave scattering length. With this definition for the two-body contact interaction, the corresponding wave function is assumed normalized to one in the $L^{2}$-norm. By giving $R$ as the length unit, with the time in units of $MR^{2}/(2\hbar)$, the dimensionless 2D GPE for the wave function $\psi\equiv\psi(\theta,\phi,t)$ is written as
@@
where $g\equiv{4\sqrt{2\pi}aN}/{\delta R}$. With $\ell$ and $m$ being the quantum numbers related to the angular variables $\phi$ and $\theta$, a given state $\psi$ is given by $\psi_{\ell m}$. By separating the $m$-angular mode, with $\psi_{\ell m}\equiv e^{{\rm i}m\phi}\Theta_{\ell m}(\theta)$, the above defined $\Delta_{\rm 2D}$ is reduced to the $1D$ operators $\Delta_{m}$, leading to the Laplace equation
@@
Existence of dark solitons — The existence of dark solitons on a bubble (dark ring solitons), previously investigated in [33, 34], can be associated with the stationary solution of the GPE,
@@
$f(\theta)=\sqrt{\frac{3}{2}}\cos\theta+\varepsilon f_{1}(\theta)+\mathcal{O}(\varepsilon^{2}),\;\;\mu=2+\frac{9}{10}\varepsilon+\mathcal{O}(\varepsilon^{2}),$ (7)
where $f_{1}(\theta)$ is uniquely defined and proportional to $P_{3}(\cos\theta)$ [see (IS9) in the SM [37]]. As the main term of (7) is the normalized $P_{1}(\cos\theta)$, the orthogonality of the Legendre polynomials imply that the normalization (4) for small $\varepsilon$ is restricted to $\mathcal{O}(\varepsilon^{2})$.
@@
With the convenient variable change $z\equiv\frac{\sqrt{\varepsilon}}{2}\left(\frac{\pi}{2}-\theta\right)$, notice that $f_{\infty}(\theta)$ is normalized to one only if $\sqrt{\varepsilon}\to\infty$. It can be incorporated into the asymptotic expansion, as shown in [37],
@@
with $f_{\infty}^{(1)}(\theta)$ being uniquely defined, given by (IS17).
In Fig. 1(a) we present profiles of the dark soliton obtained numerically, by using the shooting method described in [37]. The profile $f(\theta)$ is monotonically decreasing with the only zero at $\theta=\pi/2$ (the equator). It is close to $\sim\cos\theta$ as $\varepsilon\to 0$ according to (7) and close to $f_{\infty}(\theta)$ as $\varepsilon\to\infty$ according to (Stability of dark solitons in a bubble Bose-Einstein condensate)–(9). We confirm the asymptotic predictions by plotting $\mu$ versus $\varepsilon$ in Fig. 1(b), which shows a good agreement between the theory and numerical results.
@@
Stability of dark solitons — Linear stability analysis [40] can be performed to verify the stability of dark solitons. In this context, the well-known Bogoliubov-de Gennes (BdG) method was considered in [15]. Within this approach, small amplitude oscillations are assumed around the stationary solution (4), which is redefined as
@@
$\left\{\begin{array}[]{c}\partial_{t}u=-\Delta_{\rm 2D}v+\varepsilon f(\theta)^{2}v-\mu v,\\ -\partial_{t}v=-\Delta_{\rm 2D}u+3\varepsilon f(\theta)^{2}u-\mu u,\end{array}\right$. (13)
@@
The dark soliton is spectrally stable with respect to the $m$-angular mode if the corresponding imaginary part of $\omega$ is zero, for all eigenvalues of (15). Let us consider, for each mode $m$, the corresponding $\omega_{m}$. As we know from the general theory [40], all eigenvalues $\omega_{m}$ of the spectral stability problem (15) are real if the linear operators $L^{-}_{m}$ and $L^{+}_{m}$ are positive. These linear operators enjoy two relations for applications of the comparison principle:
@@
and
$L_{m+1}^{\pm}-L_{m}^{\pm}=\frac{2m+1}{\sin^{2}\theta}\geq 0.$ (17)
Based on (16), in [37] we demonstrate that all eigenvalues of $L_{m}^{\pm}$ are strictly positive for $m\geq 2$ and small $\varepsilon>0$, such that all eigenvalues of $L_{m}^{+}$ remain strictly positive for $m\geq 2$ and all $\varepsilon>0$. Hence, the dark soliton is stable with respect to the $m$-angular mode for small $\varepsilon$. However, the lowest eigenvalue of $L_{m}^{-}$, defined as $\omega^{-}_{m}$, can cross zero and become negative for sufficiently large $\varepsilon$, triggering the snaking instability of the dark solitons.
Next, by considering the comparative relation (17), we demonstrate in [37] that the smallest eigenvalue of $L_{m+1}^{-}$ crosses zero for values of $\varepsilon$ larger than the smallest eigenvalue of $L_{m}^{-}$ for any given $m\geq 2$, which always occurs for large values of $m$. This implies that, for every $m\geq 2$, there exists a $\varepsilon_{m}>0$ satisfying $\varepsilon_{m}<\varepsilon_{m+1}$, such that $L_{m}^{-}$ has a simple negative eigenvalue for $\varepsilon>\varepsilon_{m}$. This further clarifies that the spectral stability problem (15) provides a real unstable eigenvalue m for $\varepsilon>\varepsilon_{m}$. As shown in Table 1, the asymptotic expression, derived in [37] for $m\gg 2$,
$\varepsilon_{m}^{th}=4m(m-1),$ (18)
@@
The analytical predictions are confirmed numerically, as shown in Fig. 2 for the angular modes with $m=$ 2, 3, 4, and 5. The onset of instability in the spectral problem (15) corresponds to the zero eigenvalues of the linear operator $L_{m}^{-}$, with the imaginary part of the frequencies dynamically characterizing the instability. The dark soliton becomes unstable when $\varepsilon\gtrsim 8.37$, with $m=2$ being the initially dominant mode. The $m=2$ mode remains dominant for larger $\varepsilon$, until $\varepsilon\approx 35$ [when $\rm{Im}(\omega_{2})=\rm{Im}(\omega_{3})$], where the mode $m=3$ starts dominating the instability. As represented in Fig. 2, for a given $\varepsilon$, the largest imaginary eigenvalue characterizes the dominant unstable mode. If a given $\varepsilon$ corresponds to a dominant mode $m\geq 2$, then the dark soliton is affected by snake instability and tends to break into $m$ vortex-antivortex pairs.
@@
Figure 3: Dark-soliton dynamics, for given time instants $t$, with dominant instability modes $m=2$ (left panels, $\varepsilon=20$), $m=3$ (center panels, $\varepsilon=50$), and $m=4$ (right panels, $\varepsilon=100$), shown by the respective densities $|\psi|^{2}$. In 3D graphics, the darker the region, the less dense it is. Panels (d), (h), and (l) show the corresponding $|\psi(\frac{\pi}{2},\phi)|^{2}$ for the same instants $t$ represented in the upper 3D plots. The dashed lines are for $t=0$; the red dotted-dashed, when snake instabilities occur; and the solid-blue lines, after the breakup in vortex pairs.
Regarding the spectral stability of dark solitons for $m=1$, we demonstrate in [37] that the Eq. (15) has only real eigenvalues $\omega$, with no instability bifurcations. Finally, for $m=0$, it is also shown in [37] that there exists a single pair of eigenvalues $\omega$ of negative energy, which are smaller than all other pairs of eigenvalues $\omega$ of positive energy for small values of $\varepsilon$. These pairs can coalesce hypothetically for large values of $\varepsilon$, triggering another instability bifurcation. However, we have no supporting numerical evidence for the occurrence of such instability.
@@
The numerical approach considered in the dark-soliton dynamics is carried out first by obtaining the stationary solutions using the shooting method [49, 48]. The time evolution employs the combined spectral method together with the finite difference method introduced in [15]. For the full numerical calculations of the GPE (Stability of dark solitons in a bubble Bose-Einstein condensate), we have assumed a time-step $\delta t=10^{-5}$, with spatial grids in $\theta$ and $\phi$ directions of sizes $256\times 256$. The respective step sizes were $\delta\theta=\pi/256\approx 0.0123$ and $\delta\phi=2\pi/256\approx 0.0245$. Details of the numerical methods are given in [37].
@@
Conclusions — In the present letter, we have reported results obtained numerically and analytically for the dynamical stability of dark solitons in a single atomic BEC trapped on the surface of a rigid spherical bubble with radius $R$ and thickness $\delta R$. All the numerical and analytical results rely on a single parameter $\varepsilon$, which encompasses all information on the repulsive two-body interaction, the number of atoms, and the thickness of the bubble surface. The results of our analysis, in terms of the azimuthal excitation angular momentum modes $m$, should be useful when considering possible experimental realizations. We expect these results to impact investigations of BEC systems in spherical geometries, which are being performed aboard the international space station [2], as well as in shell condensates made with atomic mixtures [6]. Based on the spectral stability analysis, our study in the sphere shows that for the interaction parameter $\varepsilon\lesssim 8.37$, dark solitons are stable for all $m$-angular mode excitations. For $\epsilon\gtrsim 8.37$, modes with $m\geq 2$ are excited, where the dominant $m$-angular mode induces snake-like instabilities that cause the soliton to break up with the production of $m$- vortex pairs. In summary, we are highlighting the occurrence of a phenomenon that has not been previously reported when considering predictions from purely analytical studies compared with exact computational results. Our predictions for a single BEC in a sphere should provide valuable information for ongoing experimental investigations.

### v2

@@ Abstract
The stability of nonlinear waves on curved surfaces is a problem of widespread interest across physics. Here, we establish the stability criteria for dark solitons on a spherical Bose-Einstein condensate. We demonstrate a sharp instability threshold in the nonlinear parameter, beyond which solitons decay into vortex dipoles via snake instabilities. Analytically and numerically, we prove this decay is dictated by a single unstable mode for each angular momentum $m\geq 2$, which is a universal mechanism that controls the resulting vortex state. Unlike in the full three-dimensional case, where snake instabilities lead to vortex rings, a dark soliton confined to the surface of a bubble can only decay into vortex pairs.
@@
Bose-Einstein condensed gases in spherical geometries have recently attracted attention due to experiments performed aboard the International Space Station in a microgravity environment with ultra-cold gases confined in spherical and/or ellipsoidal surfaces 2020Aveline ; 2022Carollo . These experimental investigations were designed based on previous theoretical work on shell-like potentials 2001Zobay ; perrin ; 2004Zobay . Ground-based experiments also revealed shell bubbles, by exploring two species 2022Jia , or intending to observe alternative two-dimensional (2D) closed geometries Guo2022 . Such investigations have shed new light on the physics of low-dimensional quantum gases, especially in closed 2D shells. See Ref. 2025Dubessy for a recent overview of the present status and perspectives on quantum gases in bubble traps. Several interesting problems have been studied in this context, considering fundamental physics properties in shell-like structures 2023Tononi ; 2017Padavic ; 2018Padavic ; 2019Bereta ; 2019Tononi ; 2020Moller , vortex dynamics and stability 2020Padavic ; 2021Zeng ; 2021Andriati ; 2022Caracanhas , dipole interaction effects 2012Adhikari ; 2020Diniz , Berezinskii-Kosterlitz-Thouless (BKT) transition 2022Tononi , thermodynamic properties of the gas adiabatic expansion from filled sphere to hollow one 2021Rhyno , and to our main interest, some attention has been paid to the properties of condensate mixtures trapped on a bubble 2022Wolf ; 2023brito ; 2022Sturmer . Furthermore, the study of dimers on a spherical surface shows that the dimers are squeezed in the direction orthogonal to the center of mass motion, qualitatively changing its geometry, from 2D to one-dimensional (1D), leading to two-soliton motion on a bubble surface 2025Tononi .
@@
In the context of the properties of Bose-Einstein condensates (BECs) on spherical surfaces, a topic of particular interest concerns the existence, stability, and propagation of dark solitons 2016Wang . Bubble BECs are realized when atoms are confined to a thin spherical shell, creating unique curvature effects that influence the dynamics of solitons. Dark solitons are shape-localized propagating waves manifested by depressions in the condensate density. On a bubble surface, geometric effects can alter the instability dynamics of dark solitons, as compared to their propagation on flat surfaces. The curved surface is expected to affect their dynamics, propagation speed, and stability. Unlike in flat quasi-1D or quasi-2D BECs, where dark solitons are susceptible to snake instability anglin2021 , decaying into vortices, the closed topology of a bubble is expected to alter the decay pathways, with the curvature introducing geometric constraints affecting the stability dynamics. Therefore, the soliton can experience strain due to the curvature, leading to contraction or expansion depending on the interactions.
@@
Dark solitons in a flat BEC are known to oscillate and decay due to quantum fluctuations or dissipation. In quasi-1D traps, dark solitons can be metastable and decay due to transverse instabilities 1999Burger ; 2002Muryshev . Using harmonic traps, they were studied in 2000Busch ; 2000Feder , both analytically and computationally, where it was found that the large-amplitude field modulations at a frequency resonant with the energy of a dark soliton give rise to a state with multiple vortices. The stability spectrum of the dark soliton contains complex frequencies, which disappear for sufficiently small numbers of atoms or for a large transverse confinement. The relationship between these complex modes and the snake instability was investigated numerically by real-time propagation, see also  2001Dutton . In ring geometries, dark solitons in BEC systems, introduced in 1994Kivshar , were further investigated in 2003Theocharis . Finally, it is also important to note the growing interest in dark-soliton dynamics driven by the application of holographic methods to spherical superfluids, as verified in 2025Gao ; 2025Gao-pp .
@@
Two-component BECs were considered in 2001Anderson , where the dark soliton exists in one of the condensate components and the soliton nodal plane is filled with the second component. The filled solitons are stable for hundreds of milliseconds. By selectively removing the filling, one can make the soliton more susceptible to dynamical instabilities. For a condensate in a spherically symmetric potential, these instabilities cause the dark soliton to decay into stable vortex rings. By studying the oscillations and interactions of dark and dark-bright solitons, it was also shown in  2008Becker that the stability of solitons can be controlled in BECs, via confinement.
@@
The aim of the present study concerns on the spectral stability of dark solitons under discrete $m$-angular modes within a BEC confined on the surface of a bubble in the approximation of a 2D spherical hollow shell 2021Andriati . We show the occurrence of exactly one unstable mode for each $m\geq 2$, which induces a snake-like excitation that breaks the condensate into pieces, as observed in 2021Andriati . As verified numerically, an $m$-dominated unstable mode is followed by the creation of $m$ vortex-antivortex pairs, in agreement with our analytical predictions.
@@
Next, we first present the theoretical model for an atomic BEC confined to the surface of a perfect sphere, named a bubble. Dark solitons, together with the asymptotic approximations in the limit of small and large chemical potentials, are obtained analytically and numerically, supported by Supplemental Material (SM) SP . We then report the analytical results of the corresponding stability analysis, demonstrating they are in excellent agreement with numerical observations. Finally, we report outcomes of the nonlinear dynamics of the system, from which it follows that the instabilities result in the formation of vortex dipoles on the dark soliton condensate.
@@
Mathematical model — The present study is performed by assuming the condensate is trapped on the surface of a rigid spherical shell, aiming to mimic the cold-atom bubble experiments that are currently being performed in microgravity environments. In this approach, the system can be studied by considering a reduction of the three-dimensional (3D) Gross-Pitaevskii equation (GPE) to a corresponding 2D system described by spherical coordinates $\theta\in[0,\pi]$ and $\phi\in[0,2\pi]$, with fixed radius $R>0$, which is also assumed to be the space unit. In the dimensional reduction, $\delta R$ is also taken as being the radial thickness of the spherical bubble. The two-dimensional approximation is reasonable since excitations in the radial direction are inaccessible due to the large amount of energy required for them. This is true when the thickness $\delta R$ of the 3D spherical shell is very small compared to its radius $R$, that is, $\delta R\ll R$, as discussed in Refs.2021Andriati ; 2023brito . We also stress that our main concern is the dynamic stability of the dark soliton. We do not take into account its energetic instability, which could be triggered by a thermal cloud that is neglected here. The condensate can be described in the mean-field approach as a system of two coupled GPEs 1999Dalfovo ; 2016Stringari with the nonlinear two-body parameter $\overline{g}\equiv{4\pi\hbar^{2}{a_{s}}N}/{M}$, where $N$ is the number of atoms with mass $M$, and $a_{s}$ is the atom-atom $s$-wave scattering length, that we assume to be positive in our approach for a generally stable condensate. Given $R$ as the length unit, with the time in units of $MR^{2}/(2\hbar)$, the dimensionless 2D GPE for the wave function $\psi\equiv\psi(\theta,\phi,t)$, normalized to one, is written as
@@
where $g\equiv{4\sqrt{2\pi}a_{s}N}/{\delta R}$. With $\ell$ and $m$ being the quantum numbers related to the angular variables $\phi$ and $\theta$, a given state $\psi$ is given by $\psi_{\ell m}$. By separating the $m$-angular mode, with $\psi_{\ell m}\equiv e^{{\rm i}m\phi}\Theta_{\ell m}(\theta)$, the above defined $\Delta_{\rm 2D}$ is reduced to the $1D$ operators $\Delta_{m}$, leading to the Laplace equation
@@
Existence of dark solitons — The existence of dark solitons on a bubble (dark ring solitons), previously investigated in 1994Kivshar ; 2003Theocharis , can be associated with the stationary solution of the GPE,
@@
$f(\theta)=\sqrt{\frac{3}{2}}\cos\theta+\varepsilon f_{1}(\theta)+\mathcal{O}(\varepsilon^{2}),\;\;\mu=2+\frac{9}{10}\varepsilon+\mathcal{O}(\varepsilon^{2}).$ (7)
As shown in SP , $f_{1}(\theta)=\frac{3\sqrt{3}}{100\sqrt{2}}\left[3\cos\theta-5\cos^{3}\theta\right]$ is uniquely defined and proportional to $P_{3}(\cos\theta)$. As the main term of (7) is the normalized $P_{1}(\cos\theta)$, the orthogonality of the Legendre polynomials imply that the normalization (4) for small $\varepsilon$ is restricted to $\mathcal{O}(\varepsilon^{2})$.
@@
where $f_{\infty}(\theta)$ remains normallized to one only in the exact asymptotic limit. With the convenient variable change $z\equiv\frac{\sqrt{\varepsilon}}{2}\left(\frac{\pi}{2}-\theta\right)$, this solution can be incorporated into the asymptotic expansion. As shown in SP , where $f(\theta)\equiv g(z)$, $f_{\infty}(\theta)\equiv g_{0}(z)$, and $f_{\infty}^{(1)}(\theta)\equiv g_{1}(z)$:
@@
with $f_{\infty}^{(1)}(\theta)$ uniquely defined by $g_{1}(z)=\frac{d}{dz}[zg_{0}(z)]$.
In Fig. 1(a) we present profiles of the dark soliton obtained numerically, by using the shooting method described in SP . The profile $f(\theta)$ is monotonically decreasing with the only zero at $\theta=\pi/2$ (the equator). It is close to $\sim\cos\theta$ as $\varepsilon\to 0$ according to (7) and close to $f_{\infty}(\theta)$ as $\varepsilon\to\infty$ according to (9). We confirm the asymptotic predictions by plotting $\mu$ versus $\varepsilon$ in Fig. 1(b), which shows a good agreement between the theory and numerical results.
@@
Stability of dark solitons — Linear stability analysis GP25 can be performed to verify the stability of dark solitons. In this context, the well-known Bogoliubov-de Gennes (BdG) method was considered in 2021Andriati . Within this approach, small amplitude oscillations are assumed around the stationary solution (4), which is redefined as
@@
$\left\{\begin{array}[]{c}\partial_{t}u=-\Delta_{\rm 2D}v+\varepsilon f^{2}(\theta)v-\mu v,\\ -\partial_{t}v=-\Delta_{\rm 2D}u+3\varepsilon f^{2}(\theta)u-\mu u,\end{array}\right$. (13)
@@
The dark soliton is spectrally stable with respect to the $m$-angular mode if the corresponding imaginary part of $\omega$ is zero, for all eigenvalues of (15). Let us consider, for each mode $m$, the corresponding $\omega_{m}$. As we know from the general theory GP25 , all eigenvalues $\omega_{m}$ of the spectral stability problem (15) are real if the linear operators $L^{-}_{m}$ and $L^{+}_{m}$ are positive. These linear operators enjoy two comparative relations:
@@
${\rm and\hskip 28.45274pt}$ (17)
$L_{m+1}^{\pm}-L_{m}^{\pm}=\frac{2m+1}{\sin^{2}\theta}\geq 0.$
All the eigenvalues of $L_{m}^{\pm}$ are strictly positive for $m\geq 2$ and small $\varepsilon>0$, as shown in SP . With $L_{m}^{+}>L_{m}^{-}$ for all $\theta$, which are in the closed interval $0\leq\theta\leq\pi$ (excluding $\theta={\frac{\pi}{2}}$), as well as for $L_{m+1}^{\pm}>L_{m}^{\pm}$ with all $\theta$ in the open interval $0<\theta<\pi$, the smallest eigenvalue of $L_{m}^{-}$ is smaller than the smallest one of $L_{m}^{+}$; with the smallest eigenvalue of $L_{m}^{\pm}$ being smaller than the smallest one of $L_{m+1}^{\pm}$. The same holds for the second smallest eigenvalues of the same operators in the subspace of odd functions about $\theta=\frac{\pi}{2}$. Hence, the dark soliton is stable with respect to the $m$-angular mode for small $\varepsilon$. However, the lowest eigenvalue of $L_{m}^{-}$, defined as $\omega^{-}_{m}$, can cross zero and become negative for sufficiently large $\varepsilon$, triggering the snaking instability of the dark solitons.
Next, by considering (17), it is demonstrated in SP that the smallest eigenvalue of $L_{m+1}^{-}$ crosses zero for values of $\varepsilon$ larger than the smallest eigenvalue of $L_{m}^{-}$ for any given $m\geq 2$, which always occurs for large values of $m$. This implies that, for every $m\geq 2$, there exists a $\varepsilon_{m}>0$ satisfying $\varepsilon_{m}<\varepsilon_{m+1}$, such that $L_{m}^{-}$ has a simple negative eigenvalue for $\varepsilon>\varepsilon_{m}$. This further clarifies that the spectral stability problem (15) provides a real unstable eigenvalue $\omega_{m}$ for $\varepsilon>\varepsilon_{m}$. As shown in Table 1, the asymptotic expression, derived in SP for $m\gg 2$,
$\varepsilon_{m}^{th}\approx 4m(m-1),$ (18)
@@
The analytical predictions are confirmed numerically, as shown in Fig. 2 for the angular modes with $m=$ 2, 3, 4, and 5. The onset of instability in the spectral problem (15) corresponds to the zero eigenvalues of the linear operator $L_{m}^{-}$, with the imaginary part of the frequencies dynamically characterizing the instability. The dark soliton starts becoming unstable when $\varepsilon\gtrsim 8.37$, with $m=2$ being the initially dominant mode. By defining the lower limits of each dominant modes by $\varepsilon_{m}^{(0)}$, when ${\rm Im}(\omega_{m-1})={\rm Im}(\omega_{m})$, the $m=2$ mode remains dominant for larger $\varepsilon$, until $\varepsilon_{3}^{(0)}\approx 35$ [when $\rm{Im}(\omega_{2})=\rm{Im}(\omega_{3})$], where the mode $m=3$ starts dominating the instability. Next, we have $\varepsilon_{4}^{(0)}\approx 78$ for the mode $m=4$; and $\varepsilon_{5}^{(0)}\approx 140$ for $m=5$ (already outside the range shown in Fig. 2). As represented in Fig. 2, for a given $\varepsilon$, the largest imaginary eigenvalue characterizes the dominant unstable mode. If a given $\varepsilon$ corresponds to a dominant mode $m\geq 2$, then the dark soliton is affected by snake instability and tends to break into $m$ vortex-antivortex pairs.
@@
Figure 3: (Color on-line) Dark-soliton dynamics, for given instants $t$, with dominant instability modes $m=2$ (left panels, $\varepsilon=20$), $m=3$ (center panels, $\varepsilon=50$), and $m=4$ (right panels, $\varepsilon=100$), shown by the densities $|\psi|^{2}$. In 3D graphics, the darker the region, the less dense it is. At the equator, with $t$ corresponding to the upper panels in the same column, $|\psi(\frac{\pi}{2},\phi)|^{2}$ are shown in (d), (h), and (l). The bottom dashed lines are for $t=0$; the red dotted-dashed, when snake instabilities occur; and the solid-blue lines, after the breakup in vortex pairs.
Regarding the spectral stability of dark solitons for $m=1$, we demonstrate in SP that the Eq. (15) has only real eigenvalues $\omega$, with no instability bifurcations. Finally, for $m=0$, it is also shown in SP that there exists a single pair of eigenvalues $\omega$ of negative energy, which are smaller than all other pairs of eigenvalues $\omega$ of positive energy for small values of $\varepsilon$. These pairs can coalesce hypothetically for large values of $\varepsilon$, triggering another instability bifurcation. However, we have no supporting numerical evidence for the occurrence of such instability.
@@
The numerical approach considered in the dark-soliton dynamics is carried out first by obtaining the stationary solutions using the shooting method quinney ; gammal1999 . The time evolution employs the combined spectral method together with the finite difference method introduced in 2021Andriati . For the full numerical calculations of the GPE (Stability of dark solitons in a bubble Bose-Einstein condensate), we have assumed a time-step $\delta t=10^{-5}$, with spatial grids in $\theta$ and $\phi$ directions of sizes $256\times 256$. The respective step sizes were $\delta\theta=\pi/256\approx 0.0123$ and $\delta\phi=2\pi/256\approx 0.0245$. Details of the numerical methods are given in SP .
@@
Conclusions — In the present letter, we have reported results obtained numerically and analytically for the dynamical stability of dark solitons in a single atomic BEC trapped on the surface of a rigid spherical bubble with radius $R$ and thickness $\delta R$. All the numerical and analytical results rely on a single parameter $\varepsilon$, which encompasses all information on the repulsive two-body interaction, the number of atoms, and the thickness of the bubble surface. The results of our analysis, in terms of the azimuthal excitation angular momentum modes $m$, should be useful when considering possible experimental realizations. We expect these results to impact investigations of BEC systems in spherical geometries, which are being performed aboard the international space station 2022Carollo , as well as in shell condensates made with atomic mixtures 2022Jia . Based on the spectral stability analysis, our study in the sphere shows that for the interaction parameter $\varepsilon\lesssim 8.37$, dark solitons are stable for all $m$-angular mode excitations. For $\epsilon\gtrsim 8.37$, modes with $m\geq 2$ are excited, where the dominant $m$-angular mode induces snake-like instabilities that cause the soliton to break up with the production of $m$- vortex pairs.
In closing, we draw attention to the key findings reported by Anderson et al. 2001Anderson , who observed the formation of stable vortex rings resulting from the decay of a dark soliton formed in one component of a two-component Bose–Einstein condensate confined within a spherically symmetric potential. The dark soliton produced in one component (once the other component is removed), becomes susceptible to dynamical snake instabilities which decay into stable vortex rings. Our present results show that this is not achievable within a spherical 2D framework, where the condensate is confined to a bubble skin. To form closed loops in the core—identified as vortex rings—the third dimension is necessary. Further, we are highlighting that by reducing the 3D system to a quasi-2D geometry within a bubble, the vortex rings that typically arise from the decay of dark solitons are replaced by coupled vortices localized at the surface. While this outcome may be anticipated, we present a clear and specific rule governing the snake instability decay across different instability modes, as illustrated in Fig.  3. This finding reveals an important aspect that paves the way for further exploration, particularly using experimental setups like the one reported in  2001Anderson .
In summary, for dark solitons on a sphere, we are highlighting the occurrence of a universal mechanism that controls the resulting vortex state, based on purely analytical studies compared with exact computational results. Our predictions for a single BEC in a sphere should provide valuable information for ongoing experimental investigations.